OpenStudy (anonymous):

What is the slope and y-intercept of the line represented by this equation? 5x– 2y = 6 What is the slope and y-intercept of the line represented by this equation? 5x– 2y = 6

OpenStudy (anonymous):

The slope is 5/2 and the y intercept is -3 You start with the equation 5x-2y=6 then you minus 5x from both sides to get the equation -2y=-5x+6 then you multiply or divide each side of the equation by -1 to get 2y=5x-6 then you divide each side by 2 to get y=5/2x-3 if you look at the equation you will see that the slope= 5/2 and the y intercept=-3